OpenStudy (maleena_1):

Quilt squares are cut on the diagonal to form triangular quilt pieces. The hypotenuse of the resulting triangles are 18 inches long. What is the side length?

OpenStudy (dayakar):

can u draw rough figure

jhonyy9 (jhonyy9):

do you think it in this way ? 

jhonyy9 (jhonyy9):

2x^2 =18^2 make squar-roote from both sides xsqrt2 = 18 divide both sides by sqrt2 x = 18/sqrt2 rationalize the denominator x = 18sqrt2 / 2 x = 9sqrt2